# Explainable Machine Learning-Based Security and Privacy Protection Framework for Internet of Medical Things Systems

### v2

@@ 2.2 Related Work
Initially, centralized IDS solutions utilizing ML for anomaly detection are reviewed. Subsequently, an examination is conducted on solutions employing FL for intrusion detection. Finally, detection systems based on ML, incorporating XAI, are explored.
Within the domain of IDS-based ML featuring centralized architectures, notable instances include a security solution proposed by the authors of the study [14] for detecting malware within health app platforms. Their method involves combining Particle Swarm Optimization (PSO) to select features with the most impact on classification, and AdaBoost for attack detection, termed PSO-AdaBoost. To assess their solution, tests were conducted using the NSL-KDD dataset [15], comparing it against K-Nearest Neighbors (KNN) and Naive Bayes (NB) based on metrics such as precision, accuracy, and recall. The results demonstrated the superior performance of their solution over the baseline methods, highlighting its efficacy in outperforming established benchmarks.
The research presented in [16] introduced a methodology aimed at detecting anomalies and cyberattacks within healthcare systems. Their approach involves optimal feature selection using Recursive Feature Elimination (RFE), a process that iteratively eliminates the least relevant features using Logistic Regression (LR) or XGBoost Regressor methods. Subsequently, they employ a Multi-layer Perceptron (MLP) with parameter optimization for attack detection. Before applying their method, they conducted data preprocessing, replacing NaN values with column means and utilizing label encoding to transform categorical values into numerical ones. Performance evaluation was conducted using various metrics including accuracy, precision, recall, and f1-score. Their methods were applied to IoMT datasets such as WUSTL-EHMS [17], ECU-IoHT [18], Intensive Care Unit (ICU) [19], and ToN-IoT [20]. Furthermore, they compared their results with previous studies. The findings demonstrate that the combination of XGB Regressor for feature selection and optimized parameter settings for MLP yielded the most promising results among the tested methodologies.
In the work by [21] introduced an IDS designed specifically for connected healthcare systems. Their proposed methodology involves the utilization of Stacked Sparse Autoencoder (sSAE) to reduce dimensionality and the memory required for computing the covariance matrix. Subsequently, they employ the Sliced Gated Recurrent Unit (SGRU), a parallelized version of the RNN, achieved through processing segmentation. To measure the performance of their solution, the authors employed the AWID dataset [22]. However, recognizing the imbalance in the dataset, they solved this problem by applying the Synthetic Minority Oversampling Technique (SMOTE) to augment the original dataset with synthetic data. In their comparative analysis, the authors benchmarked their solution against DNN, Random Forest (RF), Long Short Term Memory (LSTM), LA-SMOTE-GRU [23], Auto Encoder (AE), XGBoost hybrid model, and LSTM hybrid model. Notably, their results demonstrate that their proposed solution outperforms these existing models, achieving superior performance metrics while simultaneously reducing model size and processing time.
The NIDS proposed by [24] is designed for smart healthcare enterprises. The proposed approach involves utilizing a Multidimensional DL Model. This model encompasses feature extraction utilizing Convolutional Neural Network (CNN), bidirectional LSTM, and CNN-LSTM models, each extracting 100 features. These features are then concatenated and passed through fully connected layers to facilitate intrusion detection. Moreover, the authors incorporated the visualization of learning features using t-SNE to offer insights into the data’s intrinsic structure. To evaluate the efficacy of their solution, they conducted tests on various datasets, including KISTI enterprise network payload, KDDCup-99 [25], CICIDS2017 [26], and WSN-DS [27], as well as UNSW-NB15 [28]. The criteria for evaluation employed in this study included accuracy, recall, f1-score, and precision. In comparison to LR, NB, KNN, DT, RF, Support Vector Machine (SVM), and previous solutions proposed for the KISTI dataset, their solution exhibited superior performance specifically within the KISTI dataset. However, it achieved similar performance levels to other solutions across the remaining datasets.
The authors of [29] have introduced an IDS tailored for the IoMT. Their innovative methodology incorporates PSO for feature selection followed by the utilization of DNN for intrusion detection within their proposed system. In their evaluation, the authors tested their solution on a dataset encompassing both network and medical features. To comprehensively assess their approach, they compared the results of their PSO-DNN solution with various ML and DL models, including LR, KNN, DT, RF, SVM, CNN, and LSTM on WUSTL-EHMS datasets [17]. For this comparative analysis, accuracy, precision, recall, and f1-score were utilized as evaluation metrics. Their findings revealed that the DL models surpassed the performance of traditional ML models. Furthermore, the PSO-DNN solution demonstrated a superior performance relative to conventional ML models and outperformed state-of-the-art approaches in intrusion detection within the IoMT domain.
In [30], a method for feature selection in IDS within the context of the IoMT is presented. Their proposed approach integrates the Logistic Redundancy Coefficient Upweighting (LRGU) technique into the Mutual Information Feature Selection (MIFS). This integration helps estimate feature redundancy within MIFS, allowing the selection of relevant features regardless of data distribution. Termed LRGU-MIFS. Subsequently, the selected relevant features are utilized in ML models for intrusion detection. Their methodology was evaluated using the WUSTL-EHMS dataset [17]. Performance comparisons were made among SVM, LR, RF, DT, and LSTM based on accuracy metrics. Tests were conducted by incrementally increasing the number of features involved in intrusion detection, comparing the outcomes across various ML models tested. The study’s outcomes revealed that the optimal results were achieved using the top 10 features out of the 45 identified by DT. Moreover, their approach demonstrated superior performance compared to previous studies utilizing different techniques for relevant feature selection.
The work made in [31] propose utilizing DNN for cyberattack detection within the context of IoMT. To achieve this, they have suggested employing PCA followed by Grey wolf optimization (GWO) to minimize the feature set used by the DNN in detecting attacks. In their study, they compared their solution against other prominent ML algorithms, namely SVM, RF, NB, and KNN, using the benchmark intrusion detection dataset. Their investigation revealed that their proposed solution outperforms these algorithms, demonstrating superior results while also reducing learning time.
In [32], a fog-cloud architecture is proposed for intrusion detection within IoMT, utilizing ensemble learning with NB, DT, RF, and XGBoost. Tests conducted on the ToN-IoT dataset [20], compared with prior studies, demonstrate the superiority of their solution.
The IDS proposed in [33] is tailored for detecting data flow modifications within multi-cloud healthcare systems, comprising gateways, edge cloud, and a core cloud. Their proposed approach involves leveraging Deep Hierarchical Stacked Neural Networks. This methodology entails reusing layers trained at the edge cloud level and merging them at the core cloud level to create a pre-trained model. To validate their solution, they conducted testing on UNSW-BOT-IoT [34] and UNSW-NB15 [28] datasets, including one generated by the authors. A comparative analysis was performed between their model, which incorporates reusing the trained layers, and a model that does not. Their findings indicate that their solution enhances accuracy while reducing training time.
Finally, the comparative study by [35] emphasizes the impact of combining network and medical attributes for intrusion detection in IoMT. Initially, they undertook the creation and collection of both medical and network data, along with simulating a MITM attack. Subsequently, they tested various ML algorithms, including ANN, RF, KNN, and SVM, based on metrics such as AUC, accuracy, and execution time. The results indicate that the performance of ML algorithms utilizing the combination of network and medical attributes yields superior results compared to using them separately.
Continuing our exploration, attention now turns towards IDS solutions harnessing the power of FL. A noteworthy example is found in the work of [36], where the authors present a framework for storing and privacy protection of data used to detect cyber-attacks within IoMT systems. Their approach involves distributing heterogeneous data across cloud nodes with privacy protection measures. They further suggested merging data from diverse sources using the differential privacy contractive deep autoencoder. This technique facilitates data fusion while reducing dimensions and safeguarding privacy during the learning process. These processed data sets serve as inputs for cyber-attack detection, employing the quantum DNN method. Their solution underwent testing on two datasets, namely the WUSTL-EHMS [17] and ICU datasets [19], achieving an accuracy and detection rate exceeding 99%. Remarkably, their solution surpasses the outcomes of prior research efforts in this field.
In addition, there have been proposals for IDS solutions that utilize a FL approach to strengthen privacy in the medical field. For instance, the authors of [37] suggest implementing an IDS to improve the security of the IoMT. The authors suggest implementing federated transfer learning, which utilizes DNN, in order to facilitate collaborative training of the cloud and edge models. The CICIDS2017 dataset is employed for the purpose of conducting experiments [26]. A multitude of metrics are employed in the performance evaluation process, encompassing accuracy, precision, detection rate, F1-score, training time, and testing time. The proposed model is evaluated in comparison to centralized learning approaches, including Stochastic Gradient Descent (SGD), Deep Belief Network, and SVM, in order to establish a benchmark. The obtained results indicate the effectiveness of the proposed model in terms of its ability to generalize and learn incrementally, outperforming the performance of other baseline ML/DL algorithms commonly used in traditional centralized learning approaches.
Addressing the security of Medical Cyber-Physical Systems (MCPS), a critical domain storing sensitive data, the study conducted by the authors of [38] propose an IDS based on ML and employing FL to provide a robust security solution safeguarding data privacy. Their approach involves creating clusters based on historical health data, with each cluster participating in the FL process to generate a global model collaboratively. To minimize communication overhead, the authors have allocated two modes for the end devices participating in the FL process: learning mode and testing mode. The testing mode restricts the transmission of the model, effectively reducing communication overhead. The evaluation was performed on the MIMIC dataset obtained from physionet [38], incorporating simulated attacks like Denial of Service (DoS), data modification, and injection attacks. The results demonstrate a high accuracy rate and a low False Positive Rate, indicating the effectiveness of their IDS solution. Moreover, the authors showed that an increase in the number of clients did not impact training time, implying that more data can further enhance the performance of their solution.
In a different approach, the authors of [39] proposed Hierarchical FL (HFL) based on hierarchical long-term memory. Their approach involves creating local models at the level of dew-servers, utilizing data collected from various end devices within the healthcare institution. These local models are then aggregated at the cloud level to create a global model. To evaluate their solution, the authors conducted tests on two datasets: ToN-IoT [20] and NSL-KDD [15]. Prior to analysis, they applied dataset preprocessing and reduced dimensionality using Principal Component Analysis (PCA). The results demonstrate the superiority of their solution compared to Long Short-Term Memory (LSTM), RNN, and Gated Recurrent Unit (GRU) models in terms of F1-score, accuracy, and precision.
In the realm of IDS-based ML incorporating XAI, a noteworthy instance is evident in the work presented in [40], the authors proposed the use of bidirectional SRU (Simple Recurrent Unit) with skip connections as a method for detecting attacks within IoMT networks. This approach effectively mitigates issues related to vanishing gradients, enhancing both training time and performance in attack detection. They compared the results achieved using two recurrent model variants, LSTM, and GRU, leveraging the ToN-IoT dataset [20]. Furthermore, they compared their findings with prior studies and discovered that their solution outperformed in terms of accuracy, precision, recall, and F-score. Additionally, they conducted an analysis of important features using an XAI technique called Local Interpretable Model-Agnostic Explanations (LIME).
Examining the reviewed solutions reveals a notable absence of an integrated approach that effectively leverages the advantages of ML for intrusion detection, FL for privacy protection, and XAI for interpretation and explanation within IoMT. Existing works focus on ML-based centralized IDS [36, 40, 14, 16, 21, 24, 26, 29, 30, 31, 35], where a central server collects and processes all data, leading to significant privacy risks and potential bottlenecks in data handling. While some studies have utilized DL techniques to improve the detection of cyber threats, these methods often lack transparency, operating as ”black-box” models that hinder trust and interpretability. Additionally, previous research on FL-based IDS solutions has primarily emphasized distributed learning efficiency but has failed to integrate mechanisms capable of explaining the decisions made by the models [37, 38, 39]. Although FL effectively addresses privacy concerns by decentralizing data processing, it does not inherently enhance the transparency of ML models. Furthermore, most reviewed solutions either exclude XAI methods entirely or rely exclusively on local model explanations [40]. In this context, the proposed solution in this article stands out as the initial effort to combine these diverse technologies ML, FL, and XAI into a cohesive framework tailored to enhance the security and interpretability of IoMT systems.
@@ 3.2 Local Training Process
[…]
@@
$g_{t}:\text{ is the gradient.}$.
$\alpha\ :\text{ is the learning rate.}$.
$\beta_{1}\beta_{2}:\text{ are the moment parameters.}$.
@@ 3.3 FL Porcess
Input: Initial model weights $w_{0}$, number of communication rounds $R$, number of local epochs $E$, fraction of clients $Fr$, number of patients $M$, learning rate $\eta$.
@@
         $w_{(t+1)}^{k}\leftarrow$ Patient Client Update $(k,w_{t},E)$triangleright Client $k$ updates its local model using its data and sends the updated weights back to the server.
@@
     $w_{(t+1)}\leftarrow\sum_{k+1}^{k}\frac{\eta_{k}}{\eta}w_{(t+1)}^{k}$triangleright Aggregate the local model weights from all participating clients to update the global model.
@@
for $e=0$ to $e=E-1$ do $\triangleright$ Perform local training for $E$ epochs.
@@
Personal devices, assumed to be smartphones with sufficient memory capacity to store medical data from associated medical equipment and enough computational power to update ML models, participate in data communication with the server and medical equipment. This involvement is illustrated in Figure 2 and detailed in the flowchart in Figure 3, where the proposed framework comprises nine steps.
In Step 1, the security administrator initializes the global model with random weights at the hospital server. Personal devices selected for FL download the initial weights to initialize their local models. In Step 2, medical sensors send captured data to paired personal devices using short-range wireless communication. In Step 3, the local model undergoes training to update weights, utilizing a specified number of epochs as outlined in Algorithm 2, once personal devices have accumulated a sufficient amount of data. This strategy is designed to minimize the number of communication rounds needed for model convergence, thereby reducing communication costs through the optimization of bandwidth usage. In Step 4, the updated weights of local models are transmitted to the hospital server through encrypted communication protocols. Step 5 sees the hospital server aggregating these weights to update the global model by summing the received weights and dividing by the number of selected participants in a communication round, thereby creating the global model as depicated in formula (6). Step 6 involves the global server sending updated weights to personal devices, and Steps 2-6 are repeated until the weight modifications become insignificant or the model converges. Once the model achieves the desired performance in Step 7, it is deployed for intrusion detection. Step 8 involves periodic activation of XAI. Finally, In Step 9, the outcomes derived from XAI methodologies, along with the historical data pertaining to intrusion detection, are transmitted to the CISO either at predefined periodic intervals or upon request. This transmission facilitates the iterative refinement and debugging of the ML model employed for intrusion detection. The historical intrusion detection data is made accessible to users via their personnel devices, thereby augmenting their confidence in the efficacy and reliability of the detection model. Concurrently, regulatory bodies are granted access to both the XAI results and the intrusion detection history. This access ensures compliance with internationally recognized standards and verifies adherence to ethical principles, particularly within the healthcare sector.
Such a process is well-suited for the IoMT system since it continuously generates data, which enriches the ML models used for IDS.
@@
[…]

### v3

@@ 2.2 Related Work
The landscape of IDS for the IoMT is vast and can be systematically organized by its underlying architectural and methodological approaches. This section provides a critical review, first examining solutions that leverage ML and DL within a centralized framework, then exploring those that utilize FL to preserve data privacy, and finally, investigating emerging approaches that incorporate XAI for interpretability.
Most existing research uses centralized models where data are aggregated on a server for training, with a strong emphasis on feature engineering and algorithmic optimization to maximize detection metrics. A prevalent theme is the combination of sophisticated feature selection techniques with powerful classifiers. For instance, the authors of [13] propose a method termed PSO-AdaBoost, which integrates Particle Swarm Optimization for feature selection with the AdaBoost classifier. Their evaluation on the NSL-KDD dataset [14] demonstrates performance superior to traditional algorithms like K-Nearest Neighbors (KNN) and Naive Bayes (NB). Similarly, the work in [15] utilizes Recursive Feature Elimination (RFE) with either Logistic Regression or an XGBoost Regressor to identify optimal features before employing a parameter-optimized Multi-layer Perceptron (MLP). This methodology is validated on a suite of IoMT datasets, including WUSTL-EHMS [16], ECU-IoHT [17], Intensive Care Unit (ICU) [18], and ToN-IoT [19], with the XGB Regressor and MLP combination yielding particularly promising results.
To address the challenges of high-dimensional data and imbalanced classes, more complex deep learning architectures are proposed. The study in [20] introduces a framework using a Stacked Sparse Autoencoder (sSAE) to reduce dimensionality and computational memory requirements, followed by a Sliced Gated Recurrent Unit (SGRU) for classification on the AWID dataset [21]. They address class imbalance with SMOTE and show their model outperforms numerous benchmarks, including DNN, Random Forest (RF), and LSTM hybrids, while also reducing model size. Further advancing architectural complexity, the authors of [22] design a Multidimensional DL Model for smart healthcare enterprises. This model extracts features using separate CNN, bidirectional LSTM, and CNN-LSTM sub-models, concatenates these features, and uses fully connected layers for final classification. Evaluated on datasets including a KISTI enterprise network payload, KDDCup-99 [23], CICIDS2017 [24], and WSN-DS [25], AND UNSW-NB15 [26], their solution shows particularly strong results on the enterprise dataset. Furthermore, an IDS designed for multi-cloud healthcare systems is proposed in [27]. Their approach leverages a Deep Hierarchical Stacked Neural Network, reusing layers trained at the edge cloud level to create a pre-trained model at the core cloud. This method, tested on UNSW-BOT-IoT [28] and UNSW-NB15 [26] datasets, is shown to enhance accuracy while reducing training time.
Research specifically targeting the IoMT domain often highlights the superiority of DL over traditional ML. The PSO-DNN model presented in [29], tested on a dataset containing both network and medical features from WUSTL-EHMS [16], finds that DL models consistently outperform ML models like LR, KNN, DT, RF, and SVM. This advantage is attributed to DL’s ability to handle complex, high-dimensional data. The critical importance of feature selection in this context is further explored in [30], where the LRGU-MIFS technique is introduced to select relevant features irrespective of data distribution, with optimal results found using a minimal feature set on the WUSTL-EHMS dataset. The PCA-GWO-DNN method in [31] also reports outperforming SVM, RF, NB, and KNN while reducing learning time. Beyond network features, the comparative study in [32] emphatically demonstrates that combining network and medical attributes significantly enhances the performance of algorithms like ANN, RF, KNN, and SVM compared to using either data type in isolation, underscoring the unique multi-modal nature of IoMT security.
While these centralized approaches achieve high performance, they possess fundamental limitations for the IoMT domain. First, they necessitate the collection of sensitive patient data to a central server, creating a significant privacy risk and a single point of failure. Second, the increasing complexity of these models, particularly DL architectures, renders them ”black boxes,” whose decisions are opaque and unexplainable. This lack of transparency is a major barrier to trust and adoption in healthcare, where understanding the rationale behind an alert is critical. Finally, as data volume grows, the scalability of constantly centralizing all data becomes a practical bottleneck.
To directly mitigate the privacy concerns of centralization, FL is introduced to enable collaborative model training without data ever leaving its source device. Several works have tailored FL for healthcare and IoMT environments. The framework proposed in [33] employs a differential privacy contractive deep autoencoder to fuse and privatize data from distributed cloud nodes before using a quantum DNN for detection. This solution achieves accuracy and detection rates exceeding 99% on the WUSTL-EHMS and ICU datasets. The authors of [34] advocate for federated transfer learning using DNN to enable collaborative training between edge and cloud models on the CICIDS2017 dataset [24], showing their approach outperforms centralized benchmarks like SGD and Deep Belief Networks in generalization and incremental learning.
Addressing the specific security needs of critical infrastructure, the study in [35] designs an FL-based IDS for Medical Cyber-Physical Systems (MCPS). Their method involves clustering historical data from the MIMIC dataset and implementing a dual-mode (learning/testing) protocol for end devices to minimize communication overhead during federated training. The system proves highly effective against simulated DoS, modification, and injection attacks. A different architectural approach, Hierarchical FL (HFL) based on hierarchical long-term memory, is put forward in [36]. This method constructs local models at dew-servers within a healthcare institution and aggregates them at the cloud level. Tested on the ToN-IoT and NSL-KDD datasets with PCA for dimensionality reduction, their HFL model surpasses the performance of LSTM, RNN, and GRU models.
FL effectively solves the data privacy problem inherent in centralized IDS by keeping data decentralized. However, the current research focus in FL-based IDS is almost exclusively on improving learning efficiency, accuracy, and reducing communication costs. A paramount shortcoming is that these solutions do not address the ”black box” nature of the collaboratively trained global model. The decisions made by the global model remain just as opaque and unexplainable as those in a centralized system, failing to provide the interpretability needed for clinical trust and operational actionability.
The field of XAI for security is still emerging, with few applications in IoMT. A noteworthy instance is the work in [37], which proposes a bidirectional Simple Recurrent Unit (SRU) with skip connections for attack detection on the ToN-IoT dataset. While the model itself achieves high performance, the key contribution is the supplementary use of Local Interpretable Model-agnostic Explanations (LIME) to provide post-hoc, local explanations for individual predictions, offering insights into which features are most influential for a specific alert.
The application of XAI in [37] represents a positive step but is indicative of the current limitations in the field. XAI is used as an external tool to explain a black-box model, providing only local rather than global interpretability. There is a lack of work on designing intrinsically interpretable models for IDS or on integrating XAI directly into the learning pipeline of a privacy-preserving framework like FL.
A systematic analysis of the reviewed literature identifies a persistent tripartite division within the field. Prevailing methodologies demonstrate proficiency in isolated domains yet fail to achieve a comprehensive synthesis. FL architectures [34, 35, 36] effectively address critical data privacy constraints through decentralized, collaborative model training; however, they inherently lack mechanisms for global model interpretability. Conversely, XAI techniques [37] furnish post-hoc, local explicability for model predictions but are predominantly applied within centralized paradigms, neglecting integration with privacy-preserving infrastructures. This stands in contrast to conventional centralized ML/DL approaches [33, 37, 13, 15, 20, 22, 24, 29, 30, 31, 32], which achieve high predictive performance at the expense of both data confidentiality and operational transparency. Consequently, a salient research gap persists: the absence of a unified architecture that concomitantly incorporates the privacy-by-design principles of FL, the transparent decision-making afforded by XAI, and the high analytical fidelity of advanced ML/DL models. This work is architected to address this deficit by introducing a novel IDS that synergistically amalgamates these three foundational components. The proposed framework is expressly designed to augment the security, privacy, and, fundamentally, the trustworthiness and ethical accountability of IoMT ecosystems.
@@ 3.2 Local Training Process
$\text{He-Initialization}=\mathcal{N}\left(0,\sqrt{\frac{2}{n}}\right)$ (2)
@@
$g_{t}:\text{ is the gradient.}$
$\alpha\ :\text{ is the learning rate.}$
$\beta_{1}\beta_{2}:\text{ are the moment parameters.}$
@@ 3.3 FL Porcess
Input: Initial model weights $w_{0}$, number of communication rounds $R$, number of local epochs $E$, fraction of clients FrFr, number of patients $M$, learning rate $\eta$.
@@
   $w_{(t+1)}^{k}\leftarrow$ Patient Client Update $(k,w_{t},E)$ $\triangleright$ Client $k$ updates its local model using its data and sends the updated weights back to the server.
@@
  $w_{(t+1)}\leftarrow\sum_{k+1}^{k}\frac{\eta_{k}}{\eta}w_{(t+1)}^{k}$ $\triangleright$ Aggregate the local model weights from all participating clients to update the global model.
@@
for $e=0$ to e=E1e=E-1 do $\triangleright$ Perform local training for $E$ epochs.
@@
Personal devices, assumed to be smartphones with sufficient memory capacity to store medical data from associated medical equipment and enough computational power to update ML models, participate in data communication with the server and medical equipment. This involvement is illustrated in Figure 2 and detailed in the flowchart in Figure 3, where the proposed framework comprises nine steps :
  1. 1.
    Initialization & Model Download: The security administrator initializes the global model with random weights at the hospital server. Personal devices selected for FL download these initial weights to initialize their local models.
  2. 2.
    IoMT Data Acquisition: Medical sensors send captured data to paired personal devices using short-range wireless communication (e.g., Bluetooth Low Energy), ensuring low-power data transfer.
  3. 3.
    Local Training: Once personal devices have accumulated a sufficient amount of data, local models are trained for a predefined number of epochs, as outlined in Algorithm 2. This approach reduces communication costs by minimizing the number of communication rounds required for model convergence, thereby optimizing bandwidth usage.
  4. 4.
    Secure Transmission of Updates: Updated local model weights are transmitted securely to the hospital server using encrypted communication protocols.
  5. 5.
    Federated Averaging (FedAvg) Aggregation: The hospital server aggregates received model weights using the FedAvg algorithm, by summing the weights and dividing by the number of selected participants in a communication round, to update the global model, as depicted in formula (6).
  6. 6.
    Global Model Broadcast: The updated global model weights are sent back to the participating personal devices.
  7. 7.
    Model Deployment & Intrusion Detection: Steps 2 through 6 are repeated until the weight modifications become insignificant and the model converges. Once the global model achieves the desired performance, it is deployed for real-time intrusion detection across the IoMT network.
  8. 8.
    Explanation Generation: Following deployment, the XAI process is applied to the final trained model. Using the SHAP method, feature attribution values are calculated to provide a transparent, auditable record of the model’s decision-making logic for each prediction.
  9. 9.
    Centralized Audit and Reporting: The outcomes derived from XAI analysis, along with historical intrusion detection data, are transmitted to the CISO either periodically or upon request, facilitating iterative refinement and debugging of the intrusion detection model, enhancing user trust through data accessibility on personal devices, and ensuring regulatory transparency and compliance with international healthcare standards.
This FL process is well-suited to IoMT systems, which continuously generate data, enriching the models used for intrusion detection. The integrated use of FL and XAI effectively addresses the critical constraints of privacy, limited device resources, and the need for transparent decision-making in healthcare environments.
@@
$W_{\text{new}}$ $:\text{The new global model after aggregation.}$